OpenStudy (anonymous):

David kicks a soccer ball off the ground and in the air with an initial velocity of 35 feet per second. Using the formula H(t) = -16t2 + vt + s, what is the maximum height the soccer ball reaches? 17.9 feet 18.2 feet 18.7 feet 19.1 feet

OpenStudy (anonymous):

@jim_thompson5910 @jdoe0001

jimthompson5910 (jim_thompson5910):

what are v and s in that equation?

OpenStudy (anonymous):

umm I think the v is 35, i might be wrong tho

jimthompson5910 (jim_thompson5910):

v = 35 is correct what is s

OpenStudy (anonymous):

im not sure

jimthompson5910 (jim_thompson5910):

s is the initial or starting height of the ball

OpenStudy (anonymous):

which is 0?

jimthompson5910 (jim_thompson5910):

v = 35 s = 0 yep

jimthompson5910 (jim_thompson5910):

so the equation is \[\Large H(t) = -16t^2 + 35t + 0\] and that turns into \[\Large H(t) = -16t^2 + 35t\]

OpenStudy (anonymous):

ok now do I factor it?

jimthompson5910 (jim_thompson5910):

How would you find the x coordinate of the vertex (for that equation)?

jimthompson5910 (jim_thompson5910):

no you wouldn't factor

OpenStudy (anonymous):

give me a second

jimthompson5910 (jim_thompson5910):

if you're not sure, here's a page that shows an example http://hotmath.com/hotmath_help/topics/vertex-of-a-parabola.html

OpenStudy (anonymous):

I'm still a little confused

jimthompson5910 (jim_thompson5910):

does the page help at all?

OpenStudy (anonymous):

I tried finding the vertex on mathway and it said a solution could not be calculated

jimthompson5910 (jim_thompson5910):

oh you used their special calculator?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

it could be a glitch, so why not try it by hand

jimthompson5910 (jim_thompson5910):

what are 'a' and 'b' in the case of \[\Large H(t) = -16t^2 + 35t\]

OpenStudy (anonymous):

okay .. i can try

OpenStudy (anonymous):

(35/32),(1225/24)

jimthompson5910 (jim_thompson5910):

I think you meant to say the vertex is (35/32, 1225/24)

OpenStudy (anonymous):

yes lol

jimthompson5910 (jim_thompson5910):

it's close, but not quite

OpenStudy (anonymous):

is that correct?

jimthompson5910 (jim_thompson5910):

no, you got the x coordinate correct

OpenStudy (anonymous):

oh dang it

jimthompson5910 (jim_thompson5910):

plug t = 35/32 into \[\Large H(t) = -16t^2 + 35t\] and evaluate to find the y coordinate

OpenStudy (anonymous):

(0,0)???

OpenStudy (anonymous):

wait.. (35,16)

jimthompson5910 (jim_thompson5910):

I have no idea how you're getting that

OpenStudy (anonymous):

neither do i >.<

jimthompson5910 (jim_thompson5910):

\[\Large H(t) = -16t^2 + 35t\] \[\Large H\left(\frac{35}{32}\right) = -16 \left(\frac{35}{32}\right) ^2 + 35\left(\frac{35}{32}\right)\] \[\Large H\left(\frac{35}{32}\right) = ???\]

OpenStudy (anonymous):

I'm not sure..

jimthompson5910 (jim_thompson5910):

use a calculator if you need to (wolfram alpha is a good calculator to evaluate expressions like that)

OpenStudy (anonymous):

am i just evaluating it ?

OpenStudy (anonymous):

it said the result was 1..

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

it's not 1

OpenStudy (anonymous):

thats what wolfram said

OpenStudy (anonymous):

this is really confusing

jimthompson5910 (jim_thompson5910):

can you show me?

jimthompson5910 (jim_thompson5910):

take a screenshot and post what you got

OpenStudy (anonymous):

i can't screen shot on my computer

OpenStudy (anonymous):

it won't let me put it as a fraction.. thats the problem

jimthompson5910 (jim_thompson5910):

you hit the printscreen key

jimthompson5910 (jim_thompson5910):

are you typing in 35/32 to represent \(\Large \frac{35}{32}\) ?

OpenStudy (anonymous):

yes, it won't let me put it as a fraction :/

jimthompson5910 (jim_thompson5910):

what are you typing in exactly into wolfram alpha? Paste it here

OpenStudy (anonymous):

H(35/32)

jimthompson5910 (jim_thompson5910):

that's all?

jimthompson5910 (jim_thompson5910):

you were supposed to type in the right side, not the left

OpenStudy (anonymous):

there was only one box to type it, can you send me the link to wolfram?

jimthompson5910 (jim_thompson5910):

you type in -16(35/32)^2 + 35(35/32)

OpenStudy (anonymous):

it says: exact result 1225/24

OpenStudy (anonymous):

i think i got the answer

OpenStudy (anonymous):

19.1?

jimthompson5910 (jim_thompson5910):

are you sure? are you sure it's over 24?

OpenStudy (anonymous):

i mean 64

OpenStudy (anonymous):

sorry

jimthompson5910 (jim_thompson5910):

that's fine

jimthompson5910 (jim_thompson5910):

so yeah, 1225/64 = 19.140625

OpenStudy (anonymous):

yes.. now im wondering if i send another question if you could verify my answer? i just need a second opinion

jimthompson5910 (jim_thompson5910):

ok go ahead

OpenStudy (anonymous):

If 6c = 4T2 - 50, which of the following could T possibly be? T = square root of negative 6 times C plus 50 over 4 T = square root of negative 6 times C minus 50 over 4 T = square root of 6 times C minus 50 over 4 T = square root of 6 times C plus 50 over 4

OpenStudy (anonymous):

I chose the last one

jimthompson5910 (jim_thompson5910):

does the last one say \[\Large T = \sqrt{ \frac{6c+50}{4} } \] ??

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

it's correct

OpenStudy (anonymous):

okay ... Sarah wants to hang a painting in a gallery. The painting and frame must have an area of 75 square feet. The painting is 8 feet wide by 9 feet long. Which quadratic equation can be used to determine the thickness of the frame, x? x2 + 17x - 3 = 0 x2 + 17x + 75 = 0 4x2 + 34x + 75 = 0 4x2 + 34x - 3 = 0 I chose the third one

jimthompson5910 (jim_thompson5910):

that's incorrect

jimthompson5910 (jim_thompson5910):

what equation do you start off with?

OpenStudy (anonymous):

first i multiplied 9 times 8

OpenStudy (anonymous):

wait.. i think its the first one

OpenStudy (anonymous):

am i right?

jimthompson5910 (jim_thompson5910):

hint: find the area of this rectangle 

jimthompson5910 (jim_thompson5910):

it might help to break it up like so